### v1

@@ Abstract
Context:
Aims:A sample of SiO-masing late-type stars located in the inner Galaxy is analyzed with the goal of better constraining their obscuration.
Methods: This reference sample allows us to define mathematical relations between their dereddened infrared colours and the observed colours (e.g. K-[8], K-[24]).
Results:The derived equations define a property (a locus) of these late-type stars. Therefore, they enable us to derive the interstellar extinction. With estimated spectral types, it is possible to decompose the total extinction in the two components (interstellar and envelope extinction).
Conclusions:These relations are very useful to classify extremely obscured late-type stars located in the inner Galaxy. Estimating the two extinction components is performable on an individual late-type star, independently of its surrounding, and also when a few mid-infrared measurements are available.
Discourse on infrared colours and bolometric corrections of SiO masing stars in the inner Milky Way.
@@
SiO masing stars are usually of late-spectral types, later than M4-M5. Dusty circumstellar envelopes surround the central star, where maser emission from SiO, OH, and HO molecules may originate. The starlight is absorbed, reprocessed, and reemitted at infrared wavelengths by the dusty envelope. In other words, the envelope represents a deforming/transforming mirror of the original stellar energy distribution (SED). The effect is dependent on both the interstellar absorption caused by a diffuse medium along the line of sight and on the optical depth of the envelope, which can range from zero to an equivalent of about A= 100 mag. In the inner Galaxy, the interstellar extinction can likewise reach A= 30–40 mag. Due to the severe dust obscuration and the patchy interstellar extinction, the reddening, distances, and luminosities of late-type stars cannot be easily inferred.
@@
To characterize the SED of an asymptotic giant star (AGB) star, a decomposition of its total extinction into the interstellar component and the circumstellar component is required. Messineo et al. (2005) measured the total extinction (A) for a sample of masing AGB stars and broke it down into A (interstellar) and A (envelope). Their method was straightforward and relied on multi-wavelength (single-epoch) photometric measurements of the targeted AGB star, covering the near- and mid-infrared window. They also assumed that the median extinction of nearby giants could be used to approximate the interstellar extinction.
@@
Stellar colours and bolometric corrections (BC) of an AGB star vary depending on the properties of its envelope, e.g. extent, density, and mass loss rate. Here we present correlations between stellar colours, bolometric corrections, and envelope types obtained for the sample of SiO masing stars from Messineo et al. (2018). The advantage of this sample is that it covers a wide range of interstellar extinction and infrared stellar colours. We used it to elaborate on the intrinsic colours of cold stars in Messineo (2004), Messineo et al. (2005), and Messineo et al. (2012). This time we use this reference sample to develop a new technique. We establish correlations between the intrinsic infrared colours and the observed infrared colours and envelope types, assuming that we know the Galactic interstellar absorption curve. By solving a system of equations, it is, then, technically possible to decompose the extinction into the interstellar and envelope components, independently of the star surrounding. It is possible to model the BCs as a function of the intrinsic stellar colors and to determine the apparent bolometric magnitudes.
@@
This idea is especially helpful in the innermost obscured regions of the Milky Way, where it is not possible to study the entire SED, and individual measurements in the mid-infrared are often of poor quality (e.gg., because of confusion). If one knows the nature of the star, for example, because of its maser emission, one can use the established equations for the K-[24] or K-[8] colours to determine the extinction, characterize the SED of the star, and estimate the apparent bolometric magnitude.
@@
We describe the stellar sample and the available parameters in Sect. 2. Extinction-free colors are defined and the relationships between them and the dereddened colors are described in Sect. 3. The existence of these relations allows us to determine the interstellar extinction. Using an average spectral type (such as M6), one can also estimate the envelope extinction given the interstellar extinction.
Synthetic infrared colors derived from the DUSTY models are shown and compared with the observational data in Sect. 4.
The empirical estimates of extinction are compared with previously published estimates in Sect. 5.
2 The sample: The catalog of Messineo et al. (2018)
@@
The sample consists of 572 mass-losing late-type stars which were searched for SiO maser emission (Messineo et al. 2002, 2018). The sources are located at low latitudes (). The sample had been created using near- and mid-infrared photometric measurements. The collected measurements cover from 0.9 m to 24 m and come from the publicly available surveys 2MASS, MSX, WISE, and GLIMPSE111 2MASS stands for Two Micron All Sky Survey (Skrutskie et al. 2006), DENIS for Deep Near Infrared Survey of the Southern Sky (Epchtein et al. 1994), MSX for Midcourse Space Experiment (Price et al. 2001), WISE for Wide-field Infrared Survey Explorer (Wright et al. 2010), and GLIMPSE for Galactic Legacy Infrared Mid-Plane Survey Extraordinaire (Churchwell et al. 2009). (Messineo et al. 2018). Each photometric data point is from a single epoch. Several stellar parameters were estimated by Messineo et al. (2005) and Messineo et al. (2018).
@@
Total A determination. The total extinction was estimated by Messineo et al. (2005, 2018), in the K versus K or K diagram, by shifting the SiO masing stars onto the 47 Tuc giant branch along the reddening vector – which implies by assuming a late M-type (M4-M5). The values of A span from 0 to 5.5 mag.
@@
Infrared Bolometric corrections (BC). For a given photometric band, the BC represents the correction to be added to the dereddened apparent magnitude in order to estimate the stellar apparent bolometric magnitude (m), =m-m. There is a tight relation between the infrared BC values and the intrinsic stellar colours of Messineo et al. (2018). Second-order polynomial fits to the data are given in the Appendix.
2.1 Large amplitude variables
@@
The sample of Messineo et al. (e.g. 2005) is made of Mira-like stars. Average photometric measurements should ideally be used to infer average colours and luminosities due to large periodic light variations. Amplitudes of Miras in the K band can reach up to 2 mag, while, at longer wavelengths, they are substantially smaller (e.g., at 15 m they are about 50% than those in K). Nevertheless, average quantities of the entire sample generate quantities typical of the actual average value (Messineo et al. 2004), based on the random phase of each star (absence of synchronous pulsations). Using DENIS, 2MASS, MSX, and WISE photometric flags, Messineo et al. (2018) found that at least 74% of the sample is made of long-period variables (LPV).
@@
282 out of 572 stars have matches in the Gaia DR3 main catalog (Gaia Collaboration et al. 2023), 249 have a measured G-band magnitude; of those, 72% (179) are classified as variables; the G-band values range from 9 to 20 mag, there are from 10 to 62 G-band observations per star, from 2 to 60 BP-band observations, and from 7 to 65 RP-observations. So far, Gaia DR3 has released photometric data covering 34 months of observations (Eyer et al. 2023). Two scanning laws were used; the Ecliptic Pole Scanning Law (EPSL) during the first 28 days and the Nominal Scanning Law (NSL) afterward; the observational cadence depends on the period of the spin axis (6 h) and the precession around the Sun (63 d); the scanning law favours the Ecliptic poles. At the end of the six-year mission, each star is supposed to have an average of usable photometric transits.
@@
For the targets detected as Gaia variables, the differences between the highest and lowest G-band measurements range from 0.12 mag to 3.54 mag (Eyer et al. 2023). Those marked as with “NOT_AVAILABLE” information on variability are fainter (mostly from G=18 to 21 mag) with a peak at G=20 mag, and have from 0 to 43 G-band observations, from 0 to 19 BP-band observations, and from 0 to 38 RP-observations. The differences between the highest and lowest G-band magnitudes range from 0.0 mag to 0.67 mag.
A fraction of 55% of Gaia matches are variables in Gaia and also in Messineo et al. (2018) based on infrared measurements. Gaia adds 16% of new variables (not detected with infrared flags).
@@
The amplitude, or the difference between the minimum and maximum, is typically more than G=0.8 mag for Mira AGBs (e.g., Lebzelter et al. 2023; Messineo 2022). With variations greater than 0.8 mag, 74% of the Gaia variables are most likely Miras. But, only 86 of these 179 Gaia variables – whose amplitudes and periods range from 0.12 mag to 3.34 mag and their periods from 42 d to 808 d – are currently listed in the Gaia catalog of classified LPVs (Lebzelter et al. 2023). The Gaia catalog of LPVs includes only stars with a 5-95% quantile range larger than 0.1 mag in the G-band.
@@ 3 Extinction
The extinction-free colours are an observed quantity.
Messineo et al. (2012) show that the intrinsic MSX K colours,
(K, of SiO masing stars well correlate
@@
This correlation is of great advantage because it allows us to decompose the total A in the envelope component, A(env), and the interstellar component, A(int), and to obtain estimates of the stellar bolometric magnitudes. In this exercise, we revisit this idea to determine A(int) and A(env) quantitatively by taking the following steps:
Adopt an interstellar extinction curve.
Determine the Q values for stars in Messineo et al. (2005) that have A(int) values.
Verify the existence of mathematical relations between the Q values and, independently obtained, intrinsic colours.
For stars with unknown A(int), use these relations plus the observed Q values to estimate the intrinsic colours and A(int).
Using A and A(int), estimate the circumstellar extinction A(env).
Obtain constraints on the circumstellar extinction curve, as suggested by the kind referee.
@@ 3.1 Adopted interstellar extinction ratios
3.1.1 Near-infrared extinction power law
At near-infrared wavelength, the Galactic extinction curve is usually approximated with a power law, A and it appears the same in every direction. determinations in the literature range from 1.61 to 2.2. An index was measured by Rieke & Lebofsky (1985). A model with was favored in Messineo et al. (2005) based on an analysis of the 2MASS colors of late-type stars in the inner Galaxy. measurements larger than 2.0 were published in a number of works between 2006 and 2011 (e.g. Nishiyama et al. 2006; Stead & Hoare 2009). In the nuclear Disk, Fritz et al. (2011) reported a value of 2.1. For late-type stars at the Galactic center which suffer about 2 mag of reddening in band K, the increase of index from = 1.61 to = 1.9 implied a decrease of A of 0.7 mag. A change of index from = 1.9 to = 2.1 implies another decrease of about 0.35 mag. A higher decreases the extinction and makes the stars fainter.
In Messineo et al. (2005), an index of 1.9 was adopted, A/E(J-Ks)=0.537 and A/E(H-Ks)=1.493, and, ::=2.86:1.66:1.00. Using the same calculation performed in Messineo et al. (2005), for an index of 2.1, it is obtained that A/E(J-Ks)=0.455 and A/E(H-Ks)=1.311, and, ::=3.208:1.766:1.000. Within errors, the latter ratios are consistent with the empirical measurements of Wang & Chen (2019): ::=3.115:1.679:1.000. A/E(J-Ks)=0.473 and A/E(H-Ks)=1.472. At the Galactic center, Nogueras-Lara et al. (2020) measured = 1.87 0.03 and = 1.84 0.03, the power law with index= yields = 1.82 and = 1.77; while the Wang’s ratios yield = 1.85 and = 1.68. Here, a single power law with index= is adopted.
3.1.2 Mid-infrared extinction power law
@@
As in Messineo et al. (2005), the parametric extinction curve of Rosenthal et al. (2000) (Curve 2), which was obtained in the OMC-1 cloud with H observations, was used along with two other curves. Curve 3 is a modified version of the Rosenthal curve in the 3-8 m region to incorporate the ratios from H-line observations in the Galactic center region of Fritz et al. (2011). Curve 4 was derived by Gordon et al. (2021) with Spitzer spectroscopic data of early-type stars located in the direction of a diffuse medium. The stars used by Gordon et al. are located at A 5 mag. The average effective extinction values from Curve 2, Curve 3, and Curve 4 are listed in Table 1.
@@
The effective extinction ratios for the WISE and GLIMPSE filters were also determined by Wang & Chen (2019) using broad-band photometry of red clump stars, which are primary indicators of distances due to their known luminosity and colours. These are also shown in the Table 1 and seem to agree well with our predictions from the curves.
Filter | Curve 2 (Ros) | Curve 3 (Ros+Fritz) | Curve 4 (Gordon) | Wang | C (2) | C (3) | C (4) | ||
@@
() This column is based on the interstellar extinction curve obtained by Gordon et al. (2021) using stellar spectra taken with the Spitzer satellite.
() Values obtained with broad-band photometry by Wang & Chen (2019).
The C=C+C values are calculated using Eq. (2). C(2) refers to Curve 2, C(3) refers to Curve 3, and C(4) refers to Curve 4.
3.2 Definitions of the Q parameters
Based on the parameters and of Messineo et al. (2012), we can analogously calculate an interstellar extinction-free parameter Q for each filter and each interstellar extinction curve (Table 1):

### v5

@@ Abstract
Aims. We analysed a sample of SiO-masing late-type stars located in the inner Galaxy with the goal of setting better constraints on their obscuration.
Methods. This reference sample has allowed us to define mathematical relations between their dereddened infrared colours and the observed colours (e.g. Ks-[8], Ks-[24]).
Results. The derived equations define a property (the locus) of these late-type stars. Therefore, they enable us to derive the interstellar extinction. With estimated spectral types, it is possible to decompose the total extinction in the two components (interstellar and envelope extinction).
Conclusions. These relations are useful for classifying extremely obscured late-type stars located in the inner Galaxy. Estimating the two extinction components can be performed on an individual late-type star, independently of its surroundings and even when few mid-infrared measurements are available.
@@
Typically, SiO masing stars have late spectral types, namely, later than M4-M5. Dusty circumstellar envelopes surround the central star, where maser emission from SiO, OH, and H2O molecules may originate. The starlight is absorbed, reprocessed, and re-emitted at infrared wavelengths by the dusty envelope. In other words, the envelope represents a deforming and transforming mirror of the original stellar energy distribution (SED). The effect is dependent on both the interstellar absorption caused by a diffuse medium along the line of sight and on the optical depth of the envelope, which can range from zero to an equivalent of about AV= 100 mag. In the inner Galaxy, the interstellar extinction can likewise reach AV= 30–40 mag. Due to the severe dust obscuration and the patchy interstellar extinction, the reddening, distances, and luminosities of late-type stars cannot easily be inferred.
@@
To characterize the SED of an asymptotic giant branch (AGB) star, a decomposition of its total extinction into the interstellar component and the circumstellar component is required. Messineo et al. (2005) measured the total extinction (A) for a sample of masing AGB stars and broke it down into A for the interstellar medium (ISM) and A for the envelope. Their method was straightforward and relied on multi-wavelength (single-epoch) photometric measurements of the targeted AGB star, covering the near- and mid-infrared window. They assumed that the median extinction of nearby giants could be used to approximate the interstellar extinction.
@@
Stellar colours and bolometric corrections (BC) of an AGB star vary depending on the properties of its envelope, namely, the extent, density, and mass loss rate. Here, we present correlations between stellar colours, bolometric corrections, and envelope types obtained for the sample of SiO masing stars from Messineo et al. (2018). The advantage of this sample is that it covers a wide range of interstellar extinction and infrared stellar colours. We used it to elaborate on the intrinsic colours of cold stars in Messineo (2004), Messineo et al. (2005), and Messineo et al. (2012). This time, we have used this reference sample to develop a new technique. We have established correlations between the intrinsic infrared colours and the observed infrared colours and envelope types, assuming that we know the Galactic interstellar absorption curve. By solving a system of equations, it is then technically possible to decompose the extinction into the interstellar and envelope components, independently of the star’s surroundings. It is possible to model the BCs as a function of the intrinsic stellar colours and to determine the apparent bolometric magnitudes.
@@
This idea is especially helpful in the innermost obscured regions of the Milky Way, where it is not possible to study the entire SED, and individual measurements in the mid-infrared are often of poor quality (e.g., because of confusion). If the nature of the star is known, for example, because of its maser emission, we can use the established equations for the Ks-[24] or Ks-[8] colours to determine the extinction, characterize the SED of the star, and estimate the apparent bolometric magnitude.
@@
We describe the stellar sample and the available parameters in Sect. 2. Extinction-free colours are defined and the relationships between them and the dereddened colours are described in Sect. 3. The existence of these relations allows us to determine the interstellar extinction. Using an average spectral type (such as M6), we can also estimate the envelope extinction given the interstellar extinction. Synthetic infrared colours derived from the DUSTY models are shown and compared with the observational data in Sect. 4. The empirical estimates of extinction are compared with previously published estimates in Sect. 5.
2 Sample
@@
The sample consists of 572 late-type stars in the process of losing mass, which we searched to find SiO maser emission (Messineo et al. 2002, 2018). The sources are located at low latitudes (). The sample had been created using near- and mid-infrared photometric measurements. The collected measurements cover from 0.9 m to 24 m and come from the publicly available surveys 2MASS, MSX, WISE, and GLIMPSE111 2MASS stands for Two Micron All Sky Survey (Skrutskie et al. 2006), DENIS for Deep Near Infrared Survey of the Southern Sky (Epchtein et al. 1994), MSX for Midcourse Space Experiment (Price et al. 2001), WISE for Wide-field Infrared Survey Explorer (Wright et al. 2010), and GLIMPSE for Galactic Legacy Infrared Mid-Plane Survey Extraordinaire (Churchwell et al. 2009). (Messineo et al. 2018). Each photometric data point is from a single epoch. Several stellar parameters were estimated by Messineo et al. (2005) and Messineo et al. (2018).
@@
Total A determination. The total extinction was estimated by Messineo et al. (2005, 2018) in the Ks versus Ks or Ks diagram, by shifting the SiO masing stars onto the 47 Tuc giant branch along the reddening vector – which implies a late M type (M4-M5) star. The values of A span from 0 to 5.5 mag.
@@
Infrared bolometric corrections (BC). For a given photometric band, the BC represents the correction to be added to the dereddened apparent magnitude in order to estimate the stellar apparent bolometric magnitude, (mbol), =mbol-mλ. There is a tight relation between the infrared BC values and the intrinsic stellar colours of Messineo et al. (2018). Second-order polynomial fits to the data are given in the appendix.
@@
Large-amplitude variables. The sample from Messineo et al. (2005) is made up of Mira-like stars. The average photometric measurements should ideally be used to infer average colours and luminosities due to large periodic light variations. Amplitudes of Miras in the Ks band can reach up to 2 mag, whereas at longer wavelengths, they are substantially smaller (e.g. at 15 m they are about 50% than those in Ks). Nevertheless, average quantities of the entire sample generate quantities typical of the actual average value (Messineo et al. 2004), based on the random phase of each star (absence of synchronous pulsations). Using DENIS, 2MASS, MSX, and WISE photometric flags, Messineo et al. (2018) found that at least 74% of the sample is made of long-period variables (LPV).
@@
There are 282 out of 572 stars with matches in the Gaia DR3 main catalog (Gaia Collaboration et al. 2023), 249 have a measured G-band magnitude; of those, 72% (179) are classified as variables; the G-band values range from 9 to 20 mag; there are from 10 to 62 G-band observations per star, from 2 to 60 BP-band observations, and from 7 to 65 RP-observations. So far, Gaia DR3 has released photometric data covering 34 months of observations (Eyer et al. 2023). Two scanning laws were used; the ecliptic pole scanning law (EPSL) during the first 28 days and the nominal scanning law (NSL) afterward; the observational cadence depends on the period of the spin axis (6 h) and the precession around the Sun (63 d); the scanning law favours the Ecliptic poles. At the end of the six-year mission, each star is supposed to have an average of usable photometric transits.
@@
For the targets recognised as Gaia variables, the differences between the highest and lowest G-band measurements range from 0.12 mag to 3.54 mag (Eyer et al. 2023). Those marked as ‘NOT_AVAILABLE’ regarding variability are fainter (mostly from G=18 to 21 mag), with a peak at G=20 mag and have from 0 to 43 G-band observations, from 0 to 19 BP-band observations, and from 0 to 38 RP-observations. The differences between the highest and lowest G-band magnitudes range from 0.0 mag to 0.67 mag. A fraction of 55% of Gaia matches are variables in Gaia and also in Messineo et al. (2018) based on infrared measurements. Gaia adds 16% of new variables (not detected via infrared flags).
@@
The amplitude, or the difference between the minimum and maximum, is typically more than G=0.8 mag for Mira AGBs (e.g. Lebzelter et al. 2023; Messineo 2022). With variations greater than 0.8 mag, 74% of the Gaia variables are most likely Miras. However, only 86 of these 179 Gaia variables – whose amplitudes and periods range from 0.12 mag to 3.34 mag and their periods from 42 d to 808 d – are currently listed in the Gaia catalog of classified LPVs (Lebzelter et al. 2023). The Gaia catalog of LPVs includes only stars with a 5-95% quantile range larger than 0.1 mag in the G-band.
@@ 3 Extinction
The extinction-free colours can be calculated directly from the observations.
Messineo et al. (2012) showed that the intrinsic MSX Ks colours,
(Ks, of SiO masing stars are well correlated
@@
This correlation is of great advantage because it allows us to decompose the total A in the envelope component, A(env), and the interstellar component, A(int), and to obtain estimates of the stellar bolometric magnitudes. In this exercise, we revisit this idea to determine A(int) and A(env) quantitatively by taking the following steps.
By having adopted an interstellar extinction curve, we determine the Qλ values for stars in Messineo et al. (2005) that have A(int) values. Then, we verify the existence of mathematical relations between the Qλ values and independently obtained, intrinsic colours. For stars with unknown A(int), we use these relations plus the observed Qλ values to estimate the intrinsic colours and A(int). By using A and AK(int), we also estimate the circumstellar extinction, A(env). Eventually, we obtain some constraints on the circumstellar extinction curve, as suggested by the kind referee.
@@ 3.1 Adopted interstellar extinction ratios
3.1.1 NIR extinction power law
At near-infrared (NIR) wavelengths, the Galactic extinction curve is usually approximated with a power law, Aλ and it appears the same in every direction. determinations in the literature range from 1.61 to 2.2. A value of was measured by Rieke & Lebofsky (1985). A model with was favored in Messineo et al. (2005) based on an analysis of the 2MASS colours of late-type stars in the inner Galaxy. measurements larger than 2.0 were published in a number of works between 2006 and 2011 (e.g., Nishiyama et al. 2006; Stead & Hoare 2009). In the nuclear disk, Fritz et al. (2011) reported a value of 2.1. For late-type stars at the Galactic center, which suffer about 2 mag of reddening in bands Ks, the increase of from 1.61 to 1.9 implies a decrease of A of 0.7 mag. A change in from 1.9 to 2.1 implies another decrease of about 0.35 mag. A higher decreases the extinction and makes the stars fainter.
In Messineo et al. (2005), an of 1.9 was adopted, A/E(J-Ks)=0.537 and A/E(H-Ks)=1.493, and, ::=2.86:1.66:1.00. Using the same calculation performed in Messineo et al. (2005), for =2.1, it is found that A/E(J-Ks)=0.455 and A/E(H-Ks)=1.311, and, ::=3.208:1.766:1.000. Within errors, the latter ratios are consistent with the empirical measurements of Wang & Chen (2019): ::=3.115:1.679:1.000. A/E(J-Ks)=0.473 and A/E(H-Ks)=1.472. At the Galactic center, Nogueras-Lara et al. (2020) measured = 1.87 0.03 and = 1.84 0.03, the power law with = yields = 1.82 and = 1.77; while the Wang’s ratios yield = 1.85 and = 1.68. Here, a single power law with index = is adopted.
3.1.2 MIR extinction law
@@
As in Messineo et al. (2005), the parametric extinction curve of Rosenthal et al. (2000), denoted as ”curve 2,” obtained in the OMC-1 cloud with H2 observations, was used along with two other curves. Curve 3 is a modified version of the Rosenthal curve in the 3-8 m region to incorporate the ratios from H-line observations in the Galactic center region of Fritz et al. (2011).333 This is a revised version of curve 3 in Messineo et al. (2005), where the 3-8 m opacity enhancement had been taken as a constant. Curve 4 was derived by Gordon et al. (2021) with Spitzer spectroscopic data of early-type stars located in the direction of a diffuse medium. The stars used by Gordon et al. are located at AV 5 mag. The average effective extinction values from curve 2, curve 3, and curve 4 are listed in Table 1.
@@
The effective extinction ratios for the WISE and GLIMPSE filters were also determined by Wang & Chen (2019) using broad-band photometry of red clump stars, which are primary indicators of distances due to their known luminosity and colours. These are also shown in the Table 1 and seem to agree with our predictions from the curves.
Filter | Curve 2 (Ros)∗ | Curve 3 (Ros+Fritz)∗∗ | Curve 4 (Gordon)∗∗∗ | Wang∗∗∗∗ | Cλ (2) | Cλ (3) | Cλ (4) | ||
@@
(∗∗∗) This column is based on the interstellar extinction curve obtained by Gordon et al. (2021) using stellar spectra taken with the Spitzer satellite.
(∗∗∗∗) Values obtained with broad-band photometry by Wang & Chen (2019).
The Cλ=C+C values are calculated using Eq. (2) and Eq. (4). Cλ(2) refers to curve 2, Cλ(3) refers to curve 3, and Cλ(4) refers to curve 4.
3.2 Definitions of the Qλ parameters
Based on parameters and of Messineo et al. (2012), we can analogously calculate an interstellar extinction-free parameter Qλ for each filter and each interstellar extinction curve (Table 1):